Hospital A

Hospital B

Hospital C

Patient days TYD:

20,000

20,000

20,000

Gross Dept expenses:

$700,000

$700,000

$700,000

YTD

Food Cost

YTD Sales & Charges

YTD

Food Cost

YTD Sales & Charges

YTD

Food Cost

YTD Sales & Charges

Floor stock, nourishments

$15,000

$0.00

$15,000

$15,000

$15,000

$15,000

Internal catering

$30,000

$0.00

$30,000

$30,000

$30,000

$75,000

Catering cash

$10,000

$ 25,000

$10,000

$ 25,000

$10,000

$ 25,000

Cafeteria cash

$68,000

$150,000

$68,000

$150,000

$68,000

$150,000

Vending cash

$10,000

$ 20,000

$10,000

$ 20,000

$10,000

$ 20,000

Note: YTD = Year-to-Date

4. Based on the total gross department expenses at the hospitals, total food cost represents ___ % of expenses.
Answer #1

Hospital A = total food cost / gross department expenses * 100

1,33,000/7,00,000 * 100 = 19%

Hospital B = total food cost / gross department expenses * 100

1,33,000/ 7,00,000 * 100 = 19%

Hospital C = total food cost / gross department expenses * 100

1,33,000/ 7,00,000 * 100 = 19%   

total food cost of all hospitals= 1,33,000+ 1,33,000+1,33,000 = 3,99,000

gross departments expenses of all hospital = 7,00,000+7,00,000+ 7,00,000= 21,00,000

% of total food cost in total gross expenses = total food cost / gross departments expenses * 100

= 3,99,000/21,00,000 *  100 = 19%

Ans total food cost represent 19% of the expenses